OpenStudy (anonymous):

A group of college students is volunteering for homes for the community during their spring break. They are putting the finishing touches on a house they built. Working alone, Wade can paint a certain room in 4 hours. Rhonda can paint the same room in 3 hours. How long will it take them working together to paint the room? round your answer to the nearest hundredth if necessary

OpenStudy (lgbasallote):

@Lethal

OpenStudy (anonymous):

do you know this question?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Are there any answer choices?

OpenStudy (anonymous):

yea i'll upload a picutre

OpenStudy (anonymous):

Here it s

OpenStudy (anonymous):

I know the answer but I think I'd better leave this to someone that can explain. @Hero

OpenStudy (anonymous):

please tell me i really wanna pass this last section

hero (hero):

The only thing I have to give is this simple formula: \(\large \frac{WR}{W+R} = t\)

hero (hero):

W = Wade's Hours R = Rhonda's Hours

OpenStudy (anonymous):

so what would i have to plug in and where i honestly dont know math ( algebra)

OpenStudy (anonymous):

Thank you for the explanation.

hero (hero):

That wasn't an explanation. It was just a hint

OpenStudy (anonymous):

Good enough for me.

OpenStudy (lgbasallote):

"if you know what i mean"

hero (hero):

@simtran, you most likely should consult a personal tutor.

OpenStudy (anonymous):

can you please just give me the answer?

hero (hero):

Giving you the answer would not help you understand how to solve this problem on your own.

OpenStudy (anonymous):

I'll make it easier for you... \[{(4)(3) \over 4+3}=t\]

hero (hero):

Yeah that should help her out.

hero (hero):

If you can't figure it out from there, then I don't know what else to tell you.

OpenStudy (anonymous):

so 12 divided by seven?

OpenStudy (anonymous):

yes.

hero (hero):

Yup and you have to round to the nearest hundredth. Good luck.

OpenStudy (anonymous):

@lgbasallote really...

OpenStudy (anonymous):

@Lethal Ha ha ha...

hero (hero):

Did I miss the joke?

OpenStudy (anonymous):

Well lgbasallote mentioned me in here to come solve it when i wasn't on. and i always have to solve her problems so yeah.

hero (hero):

I ping people when they're not online sometimes as well.

OpenStudy (anonymous):

yup, well i am getting of for today. later skaters.